Consider the following sample data: 41 49 28 51 37 47 a. Calculate the range. b. Calculate MAD. (Round your intermediate calculations to at least 4 decimal places and final answer to 2 decimal places.) c. Calculate the sample variance. (Round your intermediate calculations to at least 4 decimal places and final answer to 2 decimal places.) d. Calculate the sample standard deviation. (Round your intermediate calculations to at least 4 decimal places and final answer to 2 decimal places.)
Answer #1

Sample = 41 49 28 51 37 47

a) Range = Maximum - Minimum = 51 - 28 = 23

b) Mean, Σχ/n = 253/6-42. 1667

x |x-ar x|
41 1.1667
49 6.8333
28 14.1667
51 8.8333
37 5.1667
47 4.8333

6.83 n. 6

c) Sample Variance, sigma^2 =

(41-42. 1667)2+49-42. 1667)24(28-42. 1667)24(51-42. 1667)24(37-42. 1667)2 +(47-42. 1667)2

75.37

d) Sample Standard deviation, σ V75.37 8.68